OpenStudy (anonymous):

Help? Rewrite with only sin x and cos x. sin 3x - cos x

OpenStudy (anonymous):

do you know how to write sin(a+b)

OpenStudy (anonymous):

That's what I thought I would do, too, @d3v13 . But these are my choicdes: A) 2 sin x - sin3x - cos x B) 2 sin x cos2x + sin x - 2 sin3x - cos x C) 2 sin3x cos4x + 1 D) 3 sin x cos x - sin3x - cos x Did I just overlook it? \[\sin 3x - \cos x\]

OpenStudy (anonymous):

any of the choices are not written only in sinx and cosx

OpenStudy (anonymous):

My bad.. A) 2 sin x - sin^3x - cos x B) 2 sin x cos^2x + sin x - 2 sin^3x - cos x C) 2 sin^3x cos^4x + 1 D) 3 sin x cos x - sin^3x - cos x

OpenStudy (anonymous):

oh makes more sense now

OpenStudy (anonymous):

Thought it would.. :)

OpenStudy (anonymous):

sin3x=sin(2x+x)=sin2xcosx + cos2xsinx

OpenStudy (anonymous):

=2sinxcos^2x + cos2xsinx

OpenStudy (anonymous):

okay got it

OpenStudy (anonymous):

=2sinx(1-sin^2x) + (1-2sin^2x)sinx =3sinx-4sin^3x

OpenStudy (anonymous):

hence we can use sin3x=3sinx-4sin^3x

OpenStudy (anonymous):

sin3x-cosx = 3sinx-4sin^3x-cosx =(2sinx-2sin^3x) + (sinx-2sin^3x-cosx)

OpenStudy (anonymous):

=[2sinx(1-sin^2x)] + (sinx-2sin^3x-cosx)

OpenStudy (anonymous):

=2sinxcos^2x + sinx - 2sin^3x - cosx answer is B

OpenStudy (anonymous):

Thank you so much! God bless you! :D